1. In a gas thermometer (constant pressure conditions), the gas has a volume of 1.00 L at 65.0 °C. The thermometer is placed in a freezer. The final volume measured is 0.250 L. What is the temperature of the freezer?

2. A mass of 1.56 g of an unknown gas exerts a pressure of 0.794 atm at 77 °C in a 1.00-L flask. What is the molar mass of the gas? (R = 0.08206 L·atm/mol·K)
1
Expert's answer
2019-10-28T07:49:31-0400

1. pV=nRT

n=pV/RT= 101.325 x 1 / 338 x 8.314= 0.036 mol.

T= pV/nR= 101.325 x 0.25 / 0.036 x 8.314 = 84.36 K.

2. pV=mRT/M

M= mRT/pV= 1.56 x 0.08206 x 350 / 0.794 x 1= 56 g/mol.